OpenStudy (anonymous):

The density of a solution prepared by dissolving 120 g urea (mol.M = 60) in 1000g of water is 1.15g/ml the molarity of the solution is?

OpenStudy (anonymous):

@Callisto @waterineyes

OpenStudy (anonymous):

i got it!!

OpenStudy (anonymous):

120g urea and 1000g water = total mass = 1,120g What is the volume of this solution: Density = mass of solution / volume of solution Volume of solution = mass of solution / density of solution Volume of solution: 1120/ 1.15 Volume of solution = 974mL You have 2 mol urea in 974mL solution Molarity = 2/0.974 = 2.05M

OpenStudy (anonymous):

great you need the formula of molarity. Morlarty=moles of solute) / (volume of solution in Liters)